Question
As many 8-cm diameter discs as possible are cut from a sheet of rectangular cardboard measuring 170 cm by 90 cm. Find the area of the sheet that is left.
Please help to solve this question.
Thank you very much.

Answer by fractalier(6550)

Well 21 discs one way and 11 disks the other way can be cut...231 disks in total...let's find thier total area and subtract it from the area of the original sheet...
A = 231 * ((pi)r^2)
A = 231 * 16pi
A = 3696*pi sq cm
The original sheet was 90 x 170 or 15300 sq cm.
Now subtract to get the leftover area
Leftover = 15300 - 3696(pi) or about 3689 sq cm